OpenStudy (anonymous):

find the domain of the function f of x equals 1 divided by the square root of 1 minus X

OpenStudy (kirbykirby):

\[ f(x)=\frac{1}{\sqrt{1-x}}\]?

OpenStudy (kirbykirby):

Since you can't divide by 0, you need the denominator to be not equal to 0, i.e. \[ \sqrt{1-x}\ne0\] Furthermore, whatever is under the square root should be positive, so you require: \[ \sqrt{1-x}>0\]. Then you just have to solve this inequality.

OpenStudy (kirbykirby):

(Notice that the requirement of \(> 0\) also encompasses the requirement of \(\ne 0\))

OpenStudy (anonymous):

That's what I thought but the test I'm studying for says the answer is X<1. Thanks for your help.

OpenStudy (kirbykirby):

yes once you simplify the inequality, you will get x < 1

OpenStudy (kirbykirby):

1 - x > 0 -x > -1 x < 1